HITRAN2008 EST ARRIVÉ

\underline{L. S. ROTHMAN, I. E. GORDON}, Harvard-Smithsonian Center for Astrophysics, Atomic and Molecular Physics Division, Cambridge MA 02138-1516, USA.

The new HITRAN molecular spectroscopic database has been released^{a}. This latest edition supersedes the previous edition of 2004 and greatly expands on the quantity and quality of spectral line parameters. The effort that went into establishing the new edition involved the participation of a vast number of spectroscopists, laboratories, and funding agencies throughout the world.

An overview of the new edition will be presented, with a focus on several molecules of prime importance for atmospheric remote-sensing experiments. Instructions for accessing the database can be found at the web site http://www.cfa.harvard.edu/hitran/

aL.S. Rothman, I.E. Gordon, A. Barbe, D.Chris Benner, P.F. Bernath, et al, “The HITRAN 2008 Molecular Spectroscopic Database,” JQSRT in press (2009).